This is a worded problem.
Two tuba players stand 100 yards apart. After each set of steps, the distance between them is one half of the previous distance. How many set of steps will they take before they are less than 10 yards apart?

Answer by checkley75(3666)

100/2=50/2=25/2=12.5/2=6.25 THUS AFTER 3 SETS OF STEPS THEY WILL BE 12.5 FEET APART & AFTER 4 SETS OF STEPS THEY WILL BE 6.25 FEET APART.